ghostmasterTy:

help me Adding fractions Finding a common denominator Add Simplify your answer if necessary 1. ⅞ = 7/8 2/4 = 4/8 11/8 = 2. 5/10 + 2/5 3. 9/12 4/6 4. ¼ = 5 /20 ⅕ = 4 /20 9/20 5. 1/3 3/9 6. 7/8 2/24 7. 1 7/8 x 2 2/6 = 8. 1 4/8 x 1 7/9 = 9. 3 3/4 x 3 3/4 = 4 x 3 =12 + 3 = 15/4 4 x 3 = 12 + 3 = 15/4 _____ 15/4 x 15/4 = 225/16 16) 225 = 14 1/16

Imagine:

Hey Ghost, I see you have placed all of the questions in this one post, could you break it down into another post, Thanks, Imagine- =)

ghostmasterTy:

ok

ghostmasterTy:

but would that make me spaming

Imagine:

No, many people break them down as it is easier for the helpers(us) to better understand what you are doing.

ghostmasterTy:

ok

ghostmasterTy:

so i close this post

Imagine:

So, the first one, what is the question asking? (feel free to use the Equation button as it is easier to see.) We will do the first one on this post.

ghostmasterTy:

1.

Imagine:

So you added: \[\frac{ 7 }{ 8 }+\frac{ 4 }{ 8 }=\frac{ 11 }{ 8 }\] Right?

Imagine:

For number 1. Am I correct on this?

ghostmasterTy:

ok

Imagine:

Okay, do you know how to turn an improper fraction into a proper fraction?

ghostmasterTy:

no

Imagine:

Okay, I will give you an example, use that to find your 11/8. EXAMPLE: \[\frac{ 24 }{ 7 }=?\] Okay, so let's see how many times 7 goes into 24. 7,14,(21)=3 times. So, we still have 3 leftover, so the final result would look like this: Proper Fraction:\[3 ~\frac{ 3 }{ 7 }\]

Imagine:

Now, can you turn an improper fraction into a proper one?